### v2

# Robust Adaptation of Foundation Models with Black-Box Visual Prompting

Abstract
With a surge of large-scale pre-trained models, parameter-efficient transfer learning (PETL) of large models has garnered significant attention. While promising, they commonly rely on two optimistic assumptions: 1) full access to the parameters of a PTM, and 2) sufficient memory capacity to cache all intermediate activations for gradient computation. However, in most real-world applications, PTMs serve as black-box APIs or proprietary software without full parameter accessibility. Besides, it is hard to meet a large memory requirement for modern PTMs. This work proposes black-box visual prompting (BlackVIP), which efficiently adapts the PTMs without knowledge of their architectures or parameters. BlackVIP has two components: 1) Coordinator and 2) simultaneous perturbation stochastic approximation with gradient correction (SPSA-GC). The Coordinator designs input-dependent visual prompts, which allow the target PTM to adapt in the wild. SPSA-GC efficiently estimates the gradient of PTM to update Coordinator. Besides, we introduce a variant, BlackVIP-SE, which significantly reduces the runtime and computational cost of BlackVIP. Extensive experiments on 19 datasets demonstrate that BlackVIPs enable robust adaptation to diverse domains and tasks with minimal memory requirements. We further provide a theoretical analysis on the generalization of visual prompting methods by presenting their connection to the certified robustness of randomized smoothing, and presenting an empirical support for improved robustness.
Index Terms:
Transfer Learning, Black-Box Optimization, Parameter-Efficient Fine-Tuning, Visual Prompting
I Introduction
Large-scale pre-trained models (PTMs), a.k.a. foundation models [1, 2, 3], have driven remarkable advances across diverse domains and applications, based on their outstanding transferability. By witnessing PTMs’ success, Parameter-Efficient Transfer Learning (PETL) methods that efficiently utilize PTMs for downstream tasks have recently emerged. While standard fine-tuning (FT) updates the entire or a large portion of parameters from a PTM, PETL methods aim to achieve a performance comparable to FT by adapting a tiny amount of learnable parameters. Among them, prompt-based approaches [4, 5, 6, 7] have been widely investigated on diverse research areas. For PTMs in the vision domain, Visual Prompt Tuning [5] is a representative example that injects a few learnable prompt tokens inside of Vision Transformer’s [8] layers and only optimizes them. In addition, text encoder-side prompt learning methods for vision-language models are also actively studied [9, 10, 11, 12]. Unlike the above approaches, Bahng et al. [6] initiate an investigation of visual prompting (VP), which attempts to learn a set of parameters in the input RGB pixel space as a visual prompt to steer the target PTM.
Figure 1: For transfer learning of large-scale pre-trained models (PTM), traditional FT updates the entire model, and VP adapts a few parameters in the input space. However, both FT and VP are only feasible if the PTM’s parameters are accessible to the public and require sufficient memory capacity for training. To relax these restrictions, we propose a black-box visual prompting that adopts a zeroth-order optimization (SPSA-GC) to bypass backpropagating a true gradient while reparameterizing the prompt with a tiny learnable network, Coordinator, that produces input-dependent prompts.
Although existing PETL methods show impressive potential to adapt PTM with only a few learnable parameters, they rely on two optimistic assumptions that limit their application scope: (1) full PTM parameter accessibility, and (2) a large memory capacity to cache all the intermediate activations for the gradient computation. However, PTMs in real-world applications are commonly served as black-box APIs and proprietary software, and they do not reveal the implementation-level information or full parameters due to commercial issues, e.g., violating model ownership. As a result, exploiting frontier-level PTMs to specific downstream tasks in a black-box setting, where we have limited accessibility to the model’s details, is a crucial yet unexplored problem. Furthermore, while PETL approaches have a few learnable parameters, they require a large memory capacity to cache intermediate activations to backpropagate the gradient. Therefore, users who want to adopt a large-scale PTM should retain sufficient memory capacity despite the tiny amount of learnable parameters.
To alleviate the above unrealistic assumptions, we pioneer black-box visual prompting (BlackVIP) approach, which enables the parameter-efficient transfer learning of pre-trained black-box vision models from the low-resource user perspective (illustrated in Figure 1). BlackVIP works based on the following two core components: 1) input-dependent visual prompting and 2) a stable zeroth-order optimization algorithm.
First, we augment an input image by attaching a learnable visual prompt per pixel. Note that input space prompting does not require accessibility to model architecture [7, 12] or the embedding layer weight [10, 11, 9]. In contrast to previous work on visual prompting [6] which manually determines a structure of the visual prompt, e.g., a small fraction of a fixed area in an image, we reparameterize the visual prompt with a tiny prompt generation network, Coordinator, that produces a corresponding visual prompt per image. As a result, Coordinator automatically designs each prompt based on the input rather than using an input-agnostic manual design. This allows BlackVIP to flexibly change the semantic meanings of an original image. By learning a reparameterized model instead of a prompt itself, we greatly reduce the number of parameters (from 69K of VP [6] to 9K or 1K), so that makes it suitable for the black-box optimization scenario.
Second, to train the Coordinator, BlackVIP adopts a zeroth-order optimization that estimates the gradient from pairs of input queries and outputs of a black-box model. This allows us to adapt any pre-trained vision model, whether its parameters are available in public or not, and can significantly reduce the required memory capacity compared to white-box tuning methods that require caching intermediate activations. For faster and more accurate optimization, we further present a new optimizer, Simultaneous Perturbation Stochastic Approximation with Gradient Correction (SPSA-GC) based on (SPSA) [13]. SPSA-GC first estimates the gradient of the target black-box model based on the output difference of perturbed parameters and then corrects the initial estimates in a momentum-based look-ahead manner. By integrating the Coordinator and SPSA-GC, BlackVIP achieves significant performance improvement over the baseline.
In addition to efficiency, ensuring robustness of PTMs to distribution shifts (caused by spurious correlations and noise) is also crucial for safe AI applications. Although previous PETL methods cannot address this directly, we show that BlackVIP can enhance the robustness of PTMs theoretically and empirically, even without any explicit robust training.
This work is an extension of a conference paper [14], which has a contribution as a leading work on black-box visual prompting. We summarize our new contributions in this paper as follows: (1) We propose a new variant of BlackVIP, BlackVIP-SE (Statistical feature-guided Efficient prompt), which is remarkably faster and computationally more efficient than the original BlackVIP while achieving similar classification accuracy with BlackVIP. (2) To investigate the broad applications of our method, we expand the scope of validation with more diverse viewpoints, such as compatibility with text prompting methods, applicability to post-training quantization models, and a white-box transfer regime. (3) For the first time, we provide a theoretical analysis on the generalization mechanisms of visual prompting methods through the lens of the certified robustness of a smoothed classifier.
II Related Works
II-A Parameter-Efficient Transfer Learning
PETL methods pursue fine-tuning of a small subset of PTM’s parameter set while achieving competitive performance compared to full FT. Some of them host new learnable modules inside of PTM [5, 15, 16] or on top of PTM representation [17, 18, 19, 20], and others pursue learning tokens of Transformer layer [21], e.g., CoOp [22] and its variants [11, 7, 23, 24]. Meanwhile, Bahng et al. [6] and its follow-up research [25, 26, 27] explore the Visual Prompting (VP) approach, which introduces a learnable prompt in the input space, neither into the embedding space nor the model’s architectural building blocks. The prompt is attached to an image (pixel-wise addition) and updated by gradient descent to optimize the downstream loss function. However, not all existing PETL methods are memory-efficient. That is, they require users to have sufficient memory capacity for backpropagation. More importantly, they assume full accessibility of PTM parameters, which can not be ensured in many real-world cases, where the model is only provided as a black-box API. Although there have been some recent attempts to perform memory-efficient fine-tuning [28, 29, 30], they still require access to full model parameters. To break these optimistic assumptions, we provide a black-box prompting approach that does not require parameter accessibility or large memory capacity.
II-B Black-Box Optimization
Frontier AI models [2] are usually provided through an API or proprietary software to protect ownership. To this end, some recent works fine-tune models via black-box optimization [31, 32, 33, 34]. These black-box tuning-based works commonly adopt zeroth-order optimization (which exploits estimated gradients) or derivative-free algorithms for parameter updates. For example, BAR [31] adopts a one-sided approximation gradient estimator. Meanwhile, BBT [32], and BBTv2 [33] adopt CMA-ES [35, 36], and RLPrompt [34] uses reinforcement learning (RL) to optimize the discrete prompts. However, It has been known that derivative-free optimizations (e.g. evolutionary algorithm) are hard to solve large-scale problems and do not guarantee convergence [37]. Besides, RL approaches are notorious for their instability and high variance [38]. This work adopts the Simultaneous Perturbation Stochastic Approximation (SPSA) [13] as a zeroth-order optimizer, which is known to efficiently provide a gradient approximation on high-dimensional problems [13, 39]. Besides, SPSA theoretically guarantees convergence, and the convergence error is linearly upper bounded by the dimension of optimization [13]. Although SPSA is designed to estimate high-dimensional gradients efficiently, we found that SPSA-based optimization still requires many queries in practice. Thus, we propose SPSA with Gradient Correction (SPSA-GC) that corrects the approximated gradients to enhance the convergence speed. This is the first work exploring the black-box optimization for large PTMs in the vision domain for general-purpose adaptation (rather than a specific domain [31]).
Figure 2: We propose an input-dependent prompt designer (Coordinator) and a new zeroth-order optimization algorithm (SPSA-GC) for Coordinator training. Both components are highlighted in the gray-shaded area. The combination of these two proposals forms a novel framework for black-box visual prompting with the following procedure: 1) Coordinator consumes an input image to generate a visual prompt that has the same shape as the image, then 2) the generated prompt is attached to the original image and fed into the black-box target model, finally 3) SPSA-GC takes two outputs induced by differently perturbed Coordinator parameters to estimate the gradient of loss function. Although BlackVIP [14] adopts another PTM as a frozen feature extractor, this extended work relaxes the reliance on another PTM with a simple projection obtained by a statistical method and names that approach BlackVIP-SE.
III Methodology
Background. We first present an outline of adversarial reprogramming and visual prompting, which originated from distinct motivations but closely related topics. Elasyed et al. [40] presented adversarial reprogramming (AR) inspired by adversarial attack [41, 42, 43]. The goal of AR is to repurpose the PTM to perform a new adversarial task. Let $x\in\mathbb{R}^{k\times k\times 3}$ be a resized image of the target adversarial task, and $\tilde{x}\in\mathbb{R}^{n\times n\times 3}$ is a zero-padded tensor that embeds $x$ in the center, where $k<n$. Given the target class of adversarial task $y_{adv}\in\{1,...,C_{tar}\}$, AR is formulated as:
$\operatorname*{arg\,min}_{W}(-\log{P(m(y_{adv})|\tilde{x}_{adv}})+||W||_{F})$
Here, $P(y|x)$ is a softmax output of PTM given $x$ and the adversarial image is constructed as $\tilde{x}_{adv}=\tilde{x}+\text{tanh}(W\odot M)$, and $W\in\mathbb{R}^{n\times n\times 3}$ is the adversarial program that is optimized, where $n$ is the image width of a pre-train dataset, $M$ is an optional mask for better visualization of the embedded target image, and $\odot$ denotes the element-wise multiplication. Given a pre-defined hard-coded mapping $m(\cdot)$ that maps labels from an adversarial task to labels of a pre-trained dataset, AR reprograms the model by learning perturbations without changing model architecture or weights. The sensitivity of neural networks to such perturbations has inspired many works that use the AR approach from transfer learning perspectives [31, 44, 45, 46] i.e., model reprogramming.
Meanwhile, motivated by the remarkable success of the prompting paradigm in NLP, Bahng et al. [6] explored a pixel space visual prompting (VP) approach for pre-trained vision models. By learning individual RGB pixels as an input space prompt, VP adapts a frozen PTM to targeted downstream tasks without any modification on the model architecture and/or weights. Given the input $x$ and corresponding label $y$, the learning objective of VP is defined as below,
$\operatorname*{arg\,min}_{\phi}-\log P(y|x+\phi)$
where $\phi$ is a learnable visual prompt and $P(y|x)$ is an output prop of PTM given $x$. During inference, VP employs the shared input-independent prompt for all images. It is noted that $\phi$ is attached to the fixed location, e.g., the outer part of an image like a frame by default. Though AR and VP use different terms and stem from distinct motivations, they share the fundamental concept: adapt a PTM to perform new tasks without modifying the model architecture or its weights. The goal of this work aligns with AR and VP, but we extend and improve them for broader applications.
We introduce our new input-dependent prompt generation module, Coordinator, with two instantiations for BlackVIP and BlackVIP-SE in Section III-A. Then, we explain the end-to-end framework with a new zeroth-order optimizer, SPSA-GC in Section III-B. Figure 2 illustrates the overall framework of our proposal.
III-A Coordinator: Prompt Reparameterization
The learning goal of prompting is to minimize the loss function of a downstream task with a frozen PTM by optimizing the input space visual prompt. Given a frozen PTM $P(y|x)$, and prompt-augmented image $\tilde{x}_{\bm{\phi}}$ corresponding to the label $y$, the training objective is formulated as:
$\operatorname*{arg\,min}_{\bm{\phi}}-\log P(y|\tilde{x}_{\bm{\phi}}).$
While VP and AR directly optimize the RGB pixels as learning parameters, we reparameterize the pixel space prompt to a neural network $v_{\bm{\phi}}(\cdot)$ parameterized by $\bm{\phi}=\{\phi_{d},\phi_{t}\}\in\mathbb{R}^{p}$. Specifically, we design a module named Coordinator composed of a fixed encoder $f(\cdot)$, which projects an input image $x$ to a low-dimensional feature $z_{x}$, followed by a lightweight decoder $g_{\phi_{d}}(\cdot)$ that generates a visual prompt from an integration of instance-specific feature $z_{x}$ and a learnable task-specific prompt trigger vector $\phi_{t}$. Consequently, the prompt-augmented image is formulated as follows:
$\tilde{x}_{\bm{\phi}}$ $=\text{clip}(x+\epsilon v_{\bm{\phi}}(x)),\quad v_{\bm{\phi}}(x)=g_{\phi_{d}}(z_{x},\phi_{t})$ (1)
where $z_{x}=f(x)$ is a feature vector of $x$ from the fixed encoder $f(\cdot)$, $\epsilon\in[0,1]$ is a hyperparameter that controls the intensity of visual prompt, and $\text{clip}(\cdot)$ denotes pixel-wise clipping operation to bound the prompted image in a valid RGB scale of target model’s input transformation [44]. We concatenate the prompt trigger vector $\phi_{t}$ with $z_{x}$ and then reshape them into a 3D feature map to feed into the decoder $g_{\phi_{d}}(\cdot)$ constructed by a stack of convolution layers. As a result, a rich semantic representation $z_{x}$ per instance and a task-specific prior $\phi_{t}$ are merged to design a valid visual prompt $v_{\bm{\phi}}(x)$ for a given image.
Unlike previous works [6, 31] that adopt a pre-defined input-independent prompt to a limited region of an image, our approach automatically designs the input-dependent prompt that covers an entire image region; thus, it has a higher capability to change the semantics of images if necessary. Thanks to this flexibility, it can cover more diverse tasks and be robust to challenging scenarios, e.g., distribution shift. We demonstrate the efficacy of our prompt design strategies in Section IV. We now provide two specifications of the encoder component $f(\cdot)$ in the following paragraphs.
III-A1 BlackVIP
We adopt an ImageNet pre-trained model with a self-supervised learning (SSL) objective as a frozen feature encoder $f(\cdot)$ by default. Although the encoder can also be a supervised counterpart or trained from scratch, we choose a frozen SSL encoder for the following three reasons: 1) It has been widely substantiated that self-supervised representation contains multiple discriminative features and spatial information [47, 48, 49, 50, 51, 52, 53], so it is more helpful to use a pre-trained SSL encoder than the label-supervised encoder to adapt PTM on various tasks. 2) ImageNet pre-trained models are currently well-publicized [54], so they can be readily adopted by local users, and do not hurt our realistic experimental setting. 3) By leveraging the frozen pre-trained encoder, we significantly reduce the number of learnable parameters compared to the case of training the encoder from scratch. The reduced dimension of optimization contributes to efficient gradient approximation.
III-A2 BlackVIP-SE
While BlackVIP allows us to enjoy the rich features as a source of prompts, it requires an additional forward pass of the frozen encoder. As a result, it increases the runtime during training iterations and inference, compared with model-free prompting baselines, which hinders broader applications. Here, we note an observation that the success of transfer learning of PTMs is not only attributed to high-level feature reuse but also learning from the low-level statistics across data instances [55]. Motivated by this finding, we hypothesize that low-level statistical features can be an effective alternative to model-based high-level semantic features for producing visual prompts to adapt the black-box target model. To that end, we devise a new variant, BlackVIP-SE (Statistical feature-guided Efficient prompt), by replacing the auxiliary PTM feature extractor with a more efficient statistical method. Specifically, we conduct the principal component analysis (PCA) directly on the flattened vectors of training images to find a new coordinate system that preserves the maximum variation of data. Then, we project each image into the PCA-induced low-dimensional space (98 by default) to get statistical features being convened to the decoder of the Coordinator [^1] We first fit the PCA on the few-shot train samples to get a projection matrix and use that matrix as a fixed projector during training and inference.. This simple modification endows two favorable merits compared to BlackVIP: 1) it remarkably reduces the runtime (See Table IV) during training and inference phases by replacing the PTM feature extractor into a single projection matrix, and 2) it also significantly reduces the number of learnable parameters by setting the projection dimension of PCA much lower than that of auxiliary PTM’s latent feature dimension in BlackVIP (See Table III) thereby decreasing the peak memory usage and computational cost.
Algorithm 1 BlackVIP and BlackVIP-SE algorithms
Downstream dataset $\mathcal{D}$, pre-trained model $P(\cdot|\cdot)$, Coordinator $v$ with encoder (or projector) $f$ and prompt decoder $g$, is parameterized by $\bm{\phi}_{i}=\{\phi_{d,i},\phi_{t,i}\}$, SPSA-GC decaying parameters $\{a_{i},c_{i}\}$, smoothness parameters $\beta$, prompt intensity $\epsilon$, and training iteration $T$.
// Initialize $\bm{\phi}_{1}=\{\phi_{d,1},\phi_{t,1}\}$, $\{a_{1},c_{1}\}$ and $m_{1}$
for $i$ in $1$ to $T$ do
     // Parse a batch $(x,y)\sim\mathcal{D}$ and design the prompt
     $v_{\bm{\phi}_{i}}(x)=g_{\phi_{d,i}}(f(x),\phi_{t,i})$
     $\tilde{x}_{\bm{\phi}}=\text{clip}(x+\epsilon v_{\bm{\phi}_{i}}(x))$
     // Draw a sample $\Delta_{i}$, set $c_{i}$, and estimate the gradient
     $L(\bm{\phi}_{i}):=-\log P_{\bm{\phi}_{i}}(y|\tilde{x}_{\bm{\phi}})$
     $\hat{e}_{i}(\bm{\phi}_{i})=\frac{L(\bm{\phi}_{i}+c_{i}\Delta_{i})-L(\bm{\phi}_{i}-c_{i}\Delta_{i})}{2c_{i}\Delta_{i}}$
     // Set $a_{i}$, and update parameters
     $m_{i+1}=\beta m_{i}-a_{i}\hat{e}_{i}(\bm{\phi}_{i}+\beta m_{i})$
     $\bm{\phi}_{i+1}=\bm{\phi}_{i}+m_{i+1}$
end for
III-B End-to-End Black-Box Visual Prompting
Unlike other PETL approaches, we consider the PTM as a black-box predictor that gives only a prediction output (i.e. post-softmax probability vector) for a given input image query. In this black-box setting, we adopt the zeroth-order optimizer, SPSA, with our deliberate modification to optimize Coordinator without requiring the oracle true gradient.
III-B1 SPSA
Spall et al. proposed Simultaneous Perturbation Stochastic Approximation (SPSA) [13, 56] that approximates the high-dimensional gradient efficiently. Given the positive decaying sequences of $a_{i}>0$ and $c_{i}\in[0,1]$, the gradient estimate, $\hat{e}$, and single-step parameter update of SPSA is described as follows:
$\hat{e}_{i}(\bm{\phi}_{i})$ $=\frac{L(\bm{\phi}_{i}+c_{i}\Delta_{i})-L(\bm{\phi}_{i}-c_{i}\Delta_{i})}{2c_{i}}\Delta_{i}^{-1}$ (2)
$\bm{\phi}_{i+1}$ $=\bm{\phi}_{i}-a_{i}\hat{e}_{i}(\bm{\phi}_{i})$ (3)
where $L$ is an objective function, $\bm{\phi}_{i}\in\mathbb{R}^{p}$ is $p$-dimensional learnable parameters, and $\Delta_{i}\in\mathbb{R}^{p}$ is a $i^{th}$-step random perturbation vector, sampled from mean-zero distributions that satisfy finite inverse momentum condition [13, 57] such as Rademacher and Segmented Uniform distribution. With only two forward evaluations, i.e., querying twice to the API service model, SPSA estimates the gradient by capturing the difference between model outputs of two different perturbation vectors, and we can use this estimated gradient to update the parameters $\bm{\phi}$ of the Coordinator.
III-B2 SPSA with Gradient Correction
Although the standard form of SPSA works well in myriad applications [58, 59, 60], it may suffer slow convergence in practice [61, 62], and this convergence speed issue gets even worse on the high-dimensional problem setting such as neural network optimization. We speculate that the source of slow convergence is its noisy gradient estimation caused by the irrelevant direction of randomly sampled perturbations or intrinsic data noise. To mitigate the estimation noise, inspired by Nesterov’s accelerated gradient (NAG) [63], we improve the parameter update rule in Eq. 3 as below:
$\bm{\phi}_{i+1}$ $=\bm{\phi}_{i}+m_{i+1}$ (4)
$m_{i+1}$ $=\beta m_{i}-a_{i}\hat{e}_{i}(\bm{\phi}_{i}+\beta m_{i})$ (5)
where $\beta\in[0,1)$ is a smoothing parameter. As clearly noted in [64], when the poor update, which increases the target loss, occurs, this NAG style update strongly pulls it back towards the original parameter $\bm{\phi}_{i}$. For example, SPSA-GC computes the direction of parameter update at position $\bm{\phi}_{i}+\beta m_{i}$ rather than $\bm{\phi}_{i}$, and obtains the momentum-based update vector $\beta m_{i}-a_{i}\hat{e}_{i}(\bm{\phi}_{i}+\beta m_{i})$, which is more conservative given the negative update. This kind of “correction” effect mitigates the high volatility of optimization trajectory, which is problematic in many cases of large-scale black-box optimization [65, 66]. Moreover, the NAG-style update rule excels in the larger learning rate optimization regime [64], which is standard practice for the VP optimization problem [6]. Given the strongly stochastic nature of SPSA and the large learning rate practice of VP, we conjecture that this gradient correction property can be highly effective in reducing variance in estimation [67] and increasing its convergence.
IV Results
TABLE I: Results on synthetic datasets: correlation shifts and varying object location scenarios.
Biased MNIST Loc-MNIST
$\rho=0.8$ $\rho=0.9$ 1:1 1:4
VP (white-box) 57.92 43.55 86.79 86.54
ZS 37.56 37.25 29.70 22.70
BAR 53.25 53.07 33.98 26.05
VP w/ SPSA-GC 60.34 53.86 16.21 25.68
BlackVIP-SE 55.73 55.93 62.92 55.57
BlackVIP 66.21 62.47 69.08 60.86
IV-A Experimental Setup
Datasets. First, to investigate the importance of prompt design, we consider two synthetic datasets that simulate the correlation shift and varying object location scenarios (see Section IV-B and Appendix). Then, we extensively evaluate our methods on 14 transfer learning benchmarks. These cover diverse visual domains and tasks, so they require understanding various visual semantics like scenes, actions, fine-grained categories, textures, satellite imagery, the number of objects, and natural objects. Moreover, to evaluate the robustness of prompt-based adaptation of PTM in the wild, we additionally consider $L_{2}$-noise perturbation scenario (in Figure 7) and three benchmark datasets of distribution shifts from WILDS [68] in the Supplementary Material.
Figure 3: Optimizer comparison for (left) loss curve and noise sensitivity analysis of 100-Dimensional Rosenbrock optimization problem and (Right) optimization of multilayer perception (MLP) network with 109K parameters and convolutional neural network (CNN) with 422K parameters on MNIST.
Backbone model and methods. We adopt CLIP ViT-B/16 [1] as a target PTM because CLIP does not require a separate classifier for different tasks and performs classification dynamically with text embedding features by text-side prompting. As the encoder part of the Coordinator, we use ImageNet pre-trained vit-mae-base checkpoint for BlackVIP and sci-kit-learn’s vanilla PCA or kernel PCA for BlackVIP-SE. As baseline methods, we consider 1) zero-shot classification (ZS) with the standard text prompt “a photo of {classname}”, 2) black-box adversarial reprogramming (BAR) [31] that embed the downsized images of the downstream task inside a learnable prompt, and 3) VP with SPSA-GC that simply replace the backpropagation in VP [6] with our SPSA-GC. Besides, we validate the effectiveness of our new prompt design on the white-box transfer learning setup, which enables users to access the true gradient of model parameters. Following the few-shot classification setting of [22], we use 16 samples per class for training and the entire test set for all evaluations by default. More details are provided in the Supplementary.
IV-B Experiments on Synthetic Datasets
IV-B1 Comparison between optimizers
We validate our SPSA-GC on the well-known optimization benchmark, the Rosenbrock function. We report the normalized loss ($\frac{|L(\theta^{*})-L(\theta)|}{|L(\theta^{*})-L(\theta_{0})|}$) where $L(\theta^{*})$ and $L(\theta_{0})$ is the loss value on the optimal and initial point, respectively, and $L(\theta)$ is a loss value on the current parameter $\theta\in\mathbb{R}^{100}$. In Fig. 3 (left), SPSA-GC shows faster and more stable convergence than Random Gradient-Free (RGF) adopted in previous works [69, 31], and even achieves a comparable result to Nesterov’s Accelerated Gradient (SGD-NAG) method, which adopts the true gradient. Besides, we simulate the noisy loss observation scenario (emulating the mini-batch optimization) by adding Gaussian noise to learning loss, i.e., $L_{noisy}(\theta)=L(\theta)+\xi$, where $\xi\sim\mathcal{(}0,scale^{2})$. As the noise increases, RGF rapidly degenerates while SPSA is still relatively stable, and our gradient correction (SPSA-GC) gives further improvement. This verifies the robust gradient approximation of SPSA-GC.
Beyond the small-scale optimization problem, we compare each method to optimize the actual overparameterized neural network models, multilayer perceptron (MLP) and convolutional neural network (CNN), on the MNIST classification dataset. We conduct hyperparameter tuning with the same budget for all optimizers (100 trials per method), regardless of the tunable hyperparameter of each optimizer. As the number of parameters reaches 109K and 422K for MLP and CNN, respectively, the RGF optimizer adopted by a previous work struggles to stably train the overparameterized models. Meanwhile, SPSA and SPSA-GC successfully optimize those models and achieve reasonable test accuracy, where SPSA-GC shows slightly better convergence speed and final accuracy with the aid of gradient correction.
IV-B2 Robustness to Correlation Shift
Next, we evaluate our method on Biased MNIST [70] to investigate the robustness of BlackVIPs’ input-dependent automatic prompt design under the feature-label correlation shift environment. Biased MNIST is a modified version of MNIST [71], constructed to validate a model’s generalization ability under color bias shift. At the train-time, each digit has a unique preassigned background color that strongly correlates with the target label. The degree of correlation is determined by the value $\rho\in[0,1]$, and the correlation ratio is reversed as $1-\rho$ at the inference-time. Results are summarized in Tab. I. In this setup, BlackVIPs remarkably outperform others (even white-box VP), and the performance gap becomes larger under the stronger correlation. This means that BlackVIPs, by flexibly modifying the semantics of the image with input-dependent visual prompts, can be beneficial in challenging adaptation scenarios wherein spurious correlation exists.
IV-B3 Robustness to Varying Object Location
For many object recognition datasets, the target objects are commonly placed in the center of the image. However, input data of foundation models in the wild may have varying object locations. We expect that BlackVIP and BlackVIP-SE adopt input-dependent prompts that cover the entire region of the image so they are still robust even if the object is not always located in the center of the image. To validate this, we create a variant of the MNIST, Loc-MNIST, by putting a real target digit on the four edges and an arbitrary fake digit in the center of the black blank image. The location of the target digit and the class of the fake digit are chosen randomly. We further consider a more challenging setup in that the fake digit is four times larger (1:4) than the real one. We summarize the results in Tab. I. Compared to the manually designed input-independent prompts (BAR and VP) that are frame-shaped by default, conditional prompts produced by BlackVIPs achieve significantly better performance, which supports the superiority of the Coordinator’s prompt design.
IV-C Few-shot Transfer Learning on Benchmarks
IV-C1 Main Results
We conduct a main validation on 14 commonly used benchmarks [22, 11, 6]. As shown in Tab. II, while BAR and VP undergo large performance variations across datasets, BlackVIPs show consistently high performance (i.e., improve performance from ZS on 13 and 14 over 14 datasets). Specifically, BAR shows promising results on tasks that require understanding coarse semantics (DTD [72], EuroSAT [73], and RESISC [74]), but does not show competitiveness in CLEVR [75] that requires visual reasoning (counting objects) by capturing the general semantics of the image. Meanwhile, BlackVIP performs well on various tasks by extending or limiting the region of attention of the black-box PTM (Fig. 4), which denotes that BlackVIP is a high-capability prompt learner that adapts a PTM to various downstream tasks. Overall, BlackVIP-SE also shows remarkable performance gains that are comparable to BlackVIP but significantly faster. We speculate that BlackVIP-SE excels in data sets in which the input feature space has a relatively low intrinsic dimensionality [76], such as SVHN, so PCA-based approaches can find a sufficiently discriminative low-dimensional manifold [77]. Meanwhile, BlackVIPs slightly underperform BAR on tasks that require scene-level understanding, which indicates that prompts attached to the entire image somewhat hinder a careful recognition of scene-level semantics, urging more exploration to address these tasks.
TABLE II: Top-1 classification accuracy on 14 benchmarks that require natural, specialized, structured, and fine-grained visual recognition. BlackVIP and BlackVIP-SE show outstanding results among input-space visual prompting methods. Win means the number of datasets that each method beats the zero-shot performance. Gray-colored values are the results of white-box prompt learning with stochastic gradient descent. All experiments are done in 16-shot training with three repeated runs.
Method Caltech Pets Cars Flowers Food Aircraft SUN DTD SVHN EuroSAT RESISC CLEVR UCF IN Avg. Win
VP (white-box) 94.2 90.2 66.9 86.9 81.8 31.8 67.1 61.9 60.4 90.8 81.4 40.8 74.2 67.4 71.1 13
ZS 92.9 89.1 65.2 71.3 86.1 24.8 62.6 44.7 18.1 47.9 57.8 14.5 66.8 66.7 57.6 -
BAR 93.8 88.6 63.0 71.2 84.5 24.5 62.4 47.0 34.9 77.2 65.3 18.7 64.2 64.6 61.4 6
VP w/ SPSA-GC 89.4 87.1 56.6 67.0 80.4 23.8 61.2 44.5 29.3 70.9 61.3 25.8 64.6 62.3 58.8 4
BlackVIP-SE 93.5 89.8 65.6 71.5 86.3 25.4 65.3 45.1 46.1 71.2 61.1 32.1 68.5 67.1 63.5 14
BlackVIP 93.7 89.7 65.6 70.6 86.6 25.0 64.7 45.2 44.3 73.1 64.5 36.8 69.1 67.1 64.0 13
Figure 4: Grad-CAM analysis on CLEVR, Pets, and UCF101.
IV-C2 Practical Usefulness
TABLE III: Train-time peak memory allocation and the number of learnable parameters on ImageNet-1K with batch size 64.
Method Peak Memory (MB) Params
ViT-B ViT-L ViT-B ViT-L
FT (white-box) 21,655 76,635 86M 304M
LP (white-box) 1,587 3,294 513K 769K
VP (white-box) 11,937 44,560 69K 69K
BAR 1,649 3,352 37K 37K
VP w/ SPSA-GC 1,665 3,369 69K 69K
BlackVIP-SE 1,319 2,540 1K 1K
BlackVIP 2,428 3,260 9K 9K
Memory and computation efficiency (Table III). First, BlackVIP and BlackVIP-SE have 9K and 1K parameters in total, respectively. This allows us to remarkably reduce the computational cost compared with other white-box approaches (such as FT and LP) or black-box approaches (BAR and VP). Furthermore, all the black-box adaptation approaches significantly reduce the train-time peak memory requirement, which is burdening for white-box approaches. Among the black-box methods, BlackVIP-SE achieves the best memory-and-computation efficiency.
Figure 5: Query efficiency. (x-axis) A number of queries and cost for achieving (y-axis) corresponding performance.
Query efficiency (Figure 5). Second, BlackVIPs show outstanding query efficiency among black-box prompting methods. For instance, by sending just 10K queries with 12 USD (based on Clarifai Vision API), we can improve the performance of a zero-shot model about two $\sim$ three times better using BlackVIP and BlackVIP-SE. Given that we pay for each query to adapt black-box API models, ensuring query efficiency does matter, and our BlackVIPs offer a favorable cost-and-performance trade-off.
TABLE IV: Runtime (second) comparison between prompting methods. Training denotes the sum of the forward prompt generation and the parameter update procedures for a single batch, and Inference indicates per batch forward time.
Runtime ZS BAR VP BlackVIP-SE BlackVIP
Training - 0.1402 0.1417 0.1233 0.2684
Inference 0.0044 0.0049 0.0049 0.0053 0.0098
Runtime analysis (Table IV). Next, we investigate the wall-clock time for the single-batch forward prompt generation and parameter update during training and inference time per batch. While BlackVIP achieves superior classification accuracy on many setups, it compromises the time complexity due to its reliance on an auxiliary feature extractor during prompt generation. This hinders the broad use of BlackVIP in some applications where the inference time is crucial. In contrast, our new variant, BlackVIP-SE, addresses this issue by replacing the feature extractor model with a single projection matrix founded by PCA. As a result, BlackVIP-SE achieves a competitive runtime compared with other baselines (BAR and VP) while significantly outperforming in terms of classification accuracy (See Table II). This reduced runtime during training and inference phases allows it to become a favorable alternative to BlackVIP in time-critical applications.
TABLE V: Combination of different black-box visual and text prompts with a frozen target model, where CuPL and LLM BBO denote large language model-based optimized prompts.
Visual Prompt Text Prompt EuroSAT DTD
- CuPL [78] 45.91 55.85
BAR CuPL [78] 55.22 55.42
BlackVIP-SE CuPL [78] 64.34 55.91
BlackVIP CuPL [78] 68.88 55.58
- LLM BBO [79] 52.80 45.15
BAR LLM BBO [79] 59.85 44.58
BlackVIP-SE LLM BBO [79] 70.05 45.41
BlackVIP LLM BBO [79] 71.80 45.59
IV-D Compatibility with black-box text prompting
Although this work focuses on prompt optimization for visual inputs, it has also been actively studied to optimize text prompts under the black-box target model assumption. As stated in the white-box setting [12], prompts for each modality differently affects the embedding space with complementary benefits. We investigate the compatibility of black-box visual prompting and two black-box text prompt optimization methods [78, 79]. Tab. V shows that our methods show strong compatibility with those text prompting methods overall.
IV-E Further Analysis: BlackVIP v.s. BlackVIP-SE
It is quite surprising to observe that PCA-based simple low-dimensional features can rival the high-dimensional features of a pre-trained visual encoder. We hypothesize that the effectiveness of BlackVIP-SE depends on the intrinsic dimension of the dataset [76], that is, as the intrinsic dimension of a dataset becomes smaller, the PCA features can be more effective in producing a meaningful visual prompt, resulting in rivaling performance with features from a pre-trained visual encoder. To verify this, we estimate the intrinsic dimensions in 13 benchmark datasets using the maximum likelihood estimation method proposed in [80, 81] and see the relationship between the relative performance of BlackVIP-SE compared to BlackVIP. As shown in Figure 6, the intrinsic dimension of the dataset is negatively correlated with the relative performance of BlackVIP-SE, which means that the effectiveness of PCA-based low-level statistic features depends on the intrinsic dimension of the dataset.
Figure 6: Relationship between intrinsic dimensionality estimates (with varying nearest neighbor parameter K of the estimator) and relative performance between BlackVIP and BlackVIP-SE across different datasets.
IV-F Broad Applications
IV-F1 White-box Transfer Learning
TABLE VI: White-box transfer learning with true gradients access. Gray rows denote methods that host a large number of parameters.
Mod. Method Params SVHN CLEVR Pets Avg.
- ZS - 18.1 14.5 89.1 40.6
- FT 86M 71.5 36.2 93.9 67.2
L &\&& V MaPLe 1.2M 66.5 40.5 92.9 66.6
L CoOp 9K 43.8 36.1 91.7 57.2
L CoCoOp 35K 48.6 29.7 93.0 57.1
V VP 69K 60.4 40.8 90.2 63.8
V VPT 73K 62.9 39.9 93.2 65.3
V Ours-SE 1K 60.5 36.1 90.0 62.2
V Ours 9K 60.4 42.6 90.8 64.6
V Ours 68K 61.8 43.6 90.5 65.3
V Ours 150K 61.4 45.4 91.4 66.1
Although the chief motivation of BlackVIP originated from the desire to break the parameter accessibility assumption and overwhelming memory requirements, our prompt reparameterization strategy not only resolves the memory issues but also improves the design of visual prompts compared to other visual prompting approaches. To validate the efficacy of the prompt designs of BlackVIPs, we replace the estimated gradients of optimizers with a true gradient computed by the first-order optimizer, assuming we have access to the entire model parameters. As shown in Table VI, our prompting approaches (corresponding to BlackVIP and BlackVIP-SE) outperform other recent prompt-based efficient tuning methods such as VP, VPT, CoOp, and CoCoOp under similar parameter capacity. Even our approach (with 150K parameters) is comparable with MaPLe [7], which requires eight times more parameters. These results imply that our innovative design of the prompt can also be leveraged in the white-box parameter-efficient tuning regime.
IV-F2 Visual Prompting for Quantized Models
TABLE VII: Image classification with quantized models (CLIP ViT-B/16). Evaluation is conducted on CPU.
Quantization Method SVHN CLEVR Inference time
16-bit, float ZS 18.0 14.7 0.3177
VP w/ SPSA-GC 34.8 21.8 0.3236
BlackVIP-SE 47.2 29.3 0.3223
BlackVIP 43.8 34.9 0.3489
8-bit, int ZS 15.0 16.6 0.0482
VP w/ SPSA-GC 14.7 14.2 0.0519
BlackVIP-SE 16.9 15.4 0.0548
BlackVIP 16.3 18.7 0.0584
While the main goal of BlackVIPs is to enable memory-efficient transfer learning without parameter accessibility, as the scale of modern state-of-the-art AI models continually skyrockets [82], reducing the inference latency is also increasingly crucial for deploying those models on diverse edge devices. Post-training quantization is a representative remedy that reduces the model size and required peak memory by converting the learned model parameters into low-bit ones. Here, we validate the visual prompting strategies on post-training quantized low-bit models. Specifically, we produce the visual prompts learned by the original 32-bit model and equip the prompts with images to feed into the converted low-bit model for performance evaluation. While the effectiveness varies across the bit-size and dataset, we confirm that (in Table VII) BlackVIPs can actually collaborate with the post-training quantization method to reduce the memory requirement. Compared with VP w/ SPSA-GC, BlackVIPs show remarkably better performance while not significantly increasing inference-time latency.
V Theoretical Understanding on the Generalization of the Visual Prompting
Figure 7: (a) Empirical verification on the normality assumption, (b) Illustration for the decision boundaries and generalization behavior of randomized smoothing (RS) and visual promptings, and (c) $L_{2}$ robustness validation. All experiments are conducted in a white-box setup using true gradient. We plot empirical distributions for the first two dimensions of visual prompts at the initial and end phases of training to confirm their asymptotic normality. Under this normality and other mild assumptions noted in Sec. V, VP and BlackVIP act as an adaptive RS classifier. The naive RS guarantees generalization to $L_{2}$ noise in an input, but it cannot change the false prediction of a base classifier. Although VP can adapt samples by learning visual prompts with labels, some samples may be falsely adapted due to VP’s input-agnostic nature, whereas BlackVIP is free from such risk. In terms of relative accuracy drop, BlackVIP shows better performance compared with the ZS classifier for severe noises while achieving far higher absolute accuracy compared with RS in less noisy setups.
Preliminary. A classifier $\mathcal{F}(\cdot)$ maps a $d$ dimensional input $x\in\mathbb{R}^{d}$ to a categorical variable $c\in\mathcal{C}$. We define a randomized smoothing (RS) classifier $\mathcal{G}(x):=\operatorname*{arg\,max}_{c}\text{Pr}(\mathcal{F}(x+\varepsilon)=c)$ that returns a class $c$ that maximizes an aggregated classification probability under Gaussian noise $\varepsilon\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$ with an intensity $\sigma$ and an identity matrix $\mathbf{I}$. Recent works [83, 84] showed that a smoothed classifier maintains its prediction against $L_{2}$-norm bounded perturbations, where the robustness guarantee of the smoothed classifier originated from the additive noise following a Gaussian distribution. Here, we note that the optimization trajectory of a visual prompt by SGD with a constant learning rate can be viewed as a sampling process from a Gaussian posterior centered at the local optimum [85, 86]. This indicates a prompted classifier imitates the behavior of a smoothed classifier, ensuring robustness against input perturbations at test time. The Gaussian approximation requires Assumptions 1-3 specified in the Appendix, which allow us to approximate the discrete-time SGD dynamics by a continuous-time stochastic process. As shown in Figure 7, where we visualize the first two dimensions of the visual prompts, although they are not Gaussian in the initial training phase, they end up approximating Gaussian distributions, showing the practicality of our assumption.
V-A Visual Prompting as Adaptive Randomized Smoothing
Recall the form of BlackVIP loss function: $L(\phi)=-\log{P(y|x+v_{\phi}(x))},$ where $P(y|x)$ is the PTM output probability for ground truth $y$ given $x$ and $v_{\phi}(x)$ is a visual prompt parameterized by $\phi$. This expression is reduced to a learning objective of VP, when $v_{\phi}(x)=\phi$. For simplicity, we assume the input-dependent prompt generation process BlackVIP as $v_{\phi}(x)\approx x\odot\phi$, the Hadamard product of input $x$ and the learnable parameter $\phi$. We optimize the visual prompt $\phi$ via SGD with a constant learning rate $\alpha$ as $\phi_{t+1}=\phi_{t}-\alpha\nabla L(\phi_{t})$. As shown in a previous work [85], given a sufficiently large time step $t$, when the $\phi_{t}$ converges to the minimum of loss function and $\nabla L(\phi_{t})$ approaches zero, $\phi^{*}=\phi_{t}=\phi_{t+1},...$ becomes samples from a stationary Gaussian as below:
$\phi^{*}\sim N(\mu_{\phi^{*}},\Sigma_{\phi^{*}}),\quad\Sigma_{\phi^{*}}\mathbf{A}+\mathbf{A}\Sigma_{\phi^{*}}=\frac{\alpha}{S}\mathbf{B}\mathbf{B}^{T},$ (6)
where $S$ is the minibatch size. We view $\phi^{*}$ as a random variable near the minimum of the loss function, showing Gaussian approximation after the time step $t$ from initialization. Here, $\mu_{\phi^{*}}:=\operatorname*{arg\,min}_{\phi}-\log{P(y|x+v_{\phi}(x))}$ is the point where our loss function is minimized, and $\mathbf{A}$ is a Hessian of loss w.r.t. the optimal point of $\phi^{*}$ and $\mathbf{B}\mathbf{B}^{T}$ is a covariance matrix of the SGD’s gradient noise (See [85]). Now, given that we repeatedly encounter the same sample data instance $x$ several times during training, we get the following statement,
$\mathcal{G}_{\phi^{*}}(x)=\operatorname*{arg\,max}_{c\in\mathcal{C}}\text{Pr}(\mathcal{F}(x+v_{\phi^{*}}(x))=c)$
[…]
where $\mathbf{D}_{x}=\textrm{diag}(x)$ given an input $x$. This statement reveals that a classifier with visual prompt learning can be interpreted as a randomized smoothing classifier $\mathcal{G}_{\phi^{*}}(x)$ where the location parameter ($\mu$) of the Gaussian distribution is adapted to a direction in which the downstream classification loss should be minimized. This allows us to leverage the robustness guarantee [83, 84] of a smoothed classifier for explaining the generalization behavior of visual prompting.
V-B Certified Robustness of Visual Prompting Classifier
In the previous section, we presented a new interpretation of visual prompting: the visual prompting classifier at the later part of the training stage is approximately an adaptive randomized smoothing classifier. In this section, we will derive a new statement on the generalization of the visual prompting classifier during the inference stage: the visual prompting classifier provides a robustness guarantee against perturbation $\delta$ within the radius $R$ for a given sample $x$. In other words, it preserves the most probable class $c$ for the sample $x+\delta$, which a smoothed classifier $\mathcal{G}_{\phi^{*}}(x)$ returns on the input $x$. Let $c_{A}$ be the most probable class for given $x$, and $p_{A}$ be the corresponding output probability of the classifier $\mathcal{F}$. The second most probable class is defined similarly as $c_{B}$ with the corresponding probability to $p_{B}$. The probabilities of $p_{A}$ and $p_{B}$ vary by given input, so it is useful to bound the range of these probabilities and explain the relationship with the smoothed classifier. We denote $\underline{p_{A}}$ as the lower bound for $p_{A}$ and $\overline{p_{B}}$ as the upper bound for $p_{B}$ as in [83]. Now, assume there exists $c_{A}\in\mathcal{C}$ and $\underline{p_{A}},\overline{p_{B}}$ such that:
$\text{Pr}(\mathcal{F}(x+v_{\phi^{*}}(x))=c_{A})&\geq\underline{p_{A}}\geq\overline{p_{B}}\geq\\ &\max_{c\neq c_{A}}\text{Pr}(\mathcal{F}(x+v_{\phi^{*}}(x))=c).\end{aligned}$
[…]
where $\text{where }R$ $=\frac{1}{2}\sqrt{\lambda_{\min}}(\Phi^{-1}(\underline{p_{A}})-\Phi^{-1}(\overline{p_{B}})).$
Here, $\Phi$ denotes the cumulative distribution function (CDF) of the standardized normal distribution and $\lambda_{\min}$ refers to the minimum eigenvalue of a covariance matrix $\textrm{Cov}(v_{\phi^{*}}(x))$, and $\sqrt{\lambda_{\min}}$ is bounded from above by $\min\{\sigma_{i}\}_{i=1}^{d}$ where $\sigma_{i}^{2}$ is the $i^{th}$-diagonal component of $\textrm{Cov}(v_{\phi^{*}}(x))$ from the Rayleigh quotient, elaborated in Supplementary.
Remark 1) Generalization of Visual Prompting Classifiers. From above statement, we observed that visual prompting classifier [^2] Given a test sample, VP and BlackVIP conduct inference with a single prompt in practice, but one can pursue a more accurate approximation of the behavior of smoothed classifier by storing SGD samples from the few last iterations and averaging the classifier’s predictions over them. $\mathcal{G}_{\phi^{*}}$ is robust around $x+v_{\phi^{*}}(x)$ within the radius $R$, which is impacted by the minimum variance of visual prompt $\min\{\sigma_{i}\}_{i=1}^{d}$ and the difference between $\underline{p_{A}}$ and $\overline{p_{B}}$. Importantly, this robustness guarantee is distinctive from that of randomized smoothing classifier $\mathcal{G}$, which ensures the robust prediction around $x$. That is, $\mathcal{G}_{\phi^{*}}$ first move the given sample $x$ with a learned offset direction $v_{\phi^{*}}(x)$ and holds the prediction of base classifier $\mathcal{F}(x+v_{\phi^{*}}(x))$ for the prompt-augmented sample $x+v_{\phi^{*}}(x)$ against a perturbation $\delta$, so that it can fix the incorrect prediction of base classifier $\mathcal{F}(x)$ by a prompt $v_{\phi}(x)$ learned to minimize the classification loss. Then, if we assume the test sample is a perturbed version of a train sample, i.e., $x_{\text{test}}=x_{\text{train}}+\delta$, and both have the same ground truth class label $c$, our $\mathcal{G}_{\phi^{*}}$ ensure generalization of the test sample within $||\delta||_{2}<R$ as it do on the unperturbed train sample, i.e., $\mathcal{G}_{\phi^{*}}(x_{\text{train}})=\mathcal{G}_{\phi^{*}}(x_{\text{test}})=c$. Here, the radius to ensure such generalization is determined by how large the minimum variance of a learned visual prompt is and how large the minimum gap between the top two confidence of the base classifier ($\underline{p_{A}}-\overline{p_{B}}$) given input $x$.
Remark 2) Comparison between VP and BlackVIP. According to the above analysis, the following three components are crucial to induce successful generalization in visual prompting classifiers: 1) learned offset $v_{\phi^{*}}(x)$, 2) minimum variance among visual prompt elements $\min\{\sigma_{i}\}_{i=1}^{d}$, and 3) the difference between output probabilities of the most and second-most probable classes from base classifier $\mathcal{F}$. Because the third component is shared for VP and BlackVIP, we focus on the first two components for comparison of VP and BlackVIP in terms of their generalization capability. Upon learning offset $v_{\phi^{*}}(x)$, VP uses a universal prompt $v_{\phi^{*}}(x)=\phi$ regardless of input $x$ while BlackVIP adopts an input-dependent prompt $v_{\phi^{*}}(x)$ which depends on the input $x$ during both the train-time and the test-time. Therefore, our BlackVIP has significantly better capability to modify the incorrect classification of the base model on train samples, $\mathcal{F}(x_{\text{train}})$, and results in generalization for test sample $\mathcal{F}(x_{\text{train}}+\delta)$. Moreover, the radius of robustness guarantee is much more flexible in BlackVIP due to its input-dependency on the variance of visual prompt. We speculate that this kind of conditional coverage of BlackVIP assists flexible generalization depending on the samples’ intrinsic variation, and results in better transfer learning performance across diverse environments.
In Figure 7, we compare the generalization behavior of RS, VP, and BlackVIP with an illustration of the classifier decision boundary (b), and empirical validation on EuroSAT 16-shot classification task under a white-box true gradient setting (c). Although RS achieves strong robustness to $L_{2}$-noise , which is expected given its theoretical property, RS cannot improve the false predictions of the base classifier. Meanwhile, VP and BlackVIP can adapt the base classifier with a few in-distribution samples, and both successfully improve its in-distribution generalization. BlackVIP shows better $L_{2}$-robustness compared to VP in practice due to its flexible prompt design principle.
VI Conclusion
We pioneer black-box visual prompting for a realistic and robust adaptation of PTMs. We propose BlackVIP and its efficient alternative, BlackVIP-SE, which reparameterizes the visual prompt with a conditional prompt generation network, Coordinator. By equipping our new optimizer, SPSA-GC, BlackVIPs do not require any access to model architecture or parameters and efficiently adapt the pre-trained vision model to various downstream tasks. Extensive empirical results on 19 datasets show that BlackVIP and BlackVIP-SE consistently achieve performance gains over the baseline on few-shot adaptation in the wilds, with minimal computational cost, memory capacity, and API queries. We further provide theoretical analyses of the generalization of visual prompting methods by providing a novel connection between the certified robustness of randomized smoothing and visual prompting.
Acknowledgments
This work was partly supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. 2021R1F1A1060117, 2022R1A4A3033874) and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2022-00143911,AI Excellence Global Innovative Leader Education Program).
Appendix: Robust Adaptation of Foundation Models with Black-Box Visual Prompting
Appendix A Details of Theoretical Analysis
A-A Normal Approximation
The approximation relating parameter updates to sampling from a normal distribution arises from the inherent stochasticity introduced by mini-batch updates in estimating the gradient of the loss function. According to the Central Limit Theorem (CLT), this gradient approaches a normal distribution as the batch size $S$ increases. The application of the CLT relies on the assumption that the stochastic gradient is the summation of independent and uniformly sampled $S$. This leads to the following convergence in distribution for the gradient noise:
[…]
where $\nabla\hat{L}_{S}(\phi)$ denotes the stochastic gradient estimate based on a mini-batch of size $S$, $\nabla{L}(\phi)$ represents the true gradient over the entire dataset. Furthermore, we assume the gradient noise, $\nabla\hat{L}_{S}(\phi)-\nabla L(\phi)$, has a covariance approximated by a constant matrix $\mathbf{C}=\mathbf{B}\mathbf{B}^{T}$, where $\mathbf{C}$ is factorized into $\mathbf{B}\mathbf{B}^{T}$.
Assumption 1 (Gaussian Gradient Noise with Constant Covariance)
In the vicinity of a local optimum, assuming a sufficiently small learning rate $\alpha$ such that iterates $\phi_{k}$ remain confined to a neighborhood, the gradient noise covariance matrix is approximated by a constant, semi-positive definite matrix $C$.
The standard discrete-time SGD update rule is given by:
$\phi_{k+1}=\phi_{k}-\alpha\nabla\hat{L}_{S}(\phi_{k})$
where $\phi_{k}$ denotes the parameter vector at iteration $k$.
Assumption 2 (Continuous-Time Approximation)
For a sufficiently small learning rate $\alpha$, the discrete-time SGD is approximated by a continuous-time stochastic differential equation(SDE).
Assumption 3 (Quadratic Loss Approximation)
Around a local minimum $\phi^{*}$, the loss function $L(\phi)$ can be approximated by a quadratic function form:
$L(\phi)\approx L(\phi^{*})+\frac{1}{2}(\phi-\phi^{*})^{T}\mathbf{A}(\phi-\phi^{*}).$
Wihtout loss of generality, the minimum is located at $\phi=0$, which simplifes the equation to
$L(\phi)\approx\frac{1}{2}\phi^{T}\mathbf{A}\phi.$
Here, we also assume positive definiteness on $\mathbf{A}$. These assumptions come from [85] and enables the continuous-time SDE of $\phi$ to reduce to a multivariate Ornstein-Uhlenbeck process, which admits a Gaussian stationary distribution, where:
$q(\phi)\propto\exp\{-\frac{1}{2}\phi^{T}\Sigma_{\phi}^{-1}\phi\}.$
$q(\phi)$ denotes a stationary distribution and the covariance $\Sigma_{\phi}$ satisfies $\Sigma_{\phi}\mathbf{A}+\mathbf{A}\Sigma_{\phi}=\frac{\alpha}{S}\mathbf{B}\mathbf{B}^{T}$, with $\mathbf{A}$ being the hessian matrix at the optimal point.
A-B Randomized Smoothing and Certified Robustness
A smoothed classifier exhibits robust properties within a radius $R$. Specifically, when an input is perturbed within the ball $B(x,R)=\{y:||x-y||<R\}$, the classification consistently predicts the most probable class. Building upon the foundational work by [83] which introduced a certified radius for a smoothed classifier using an isotropic Gaussian covariance, and further extensions by [84] to an anisotropic Gaussian model with a diagonal covariance matrix, our analysis advances these models by relaxing the diagonal constraint on the covariance matrix, thus proposing a more generalized form and redefining the radius based on the derived proof.
To set the stage for our proof, consider the following settings:
$X=x+\phi^{*}\sim\mathcal{N}_{d}(x+\mu_{\phi^{*}},\Sigma_{\phi^{*}})$
$X^{*}=x+\phi^{*}+\delta\sim\mathcal{N}_{d}(x+\mu_{\phi^{*}}+\delta,\Sigma_{\phi^{*}}).$
We examine the ratio of the probability density functions of $X^{*}$ and $X$, which manifests in an exponential form with the precision matrix $K=\Sigma_{\phi^{*}}^{-1}$. The ratio $\frac{p_{X^{*}}(z)}{p_{X}(z)}$ is given by:
$\frac{\exp\{-\frac{1}{2}(z-(x+\mu_{\phi^{*}}+\delta))^{T}K(z-(x+\mu_{\phi^{*}}+\delta))\}}{\exp\{-\frac{1}{2}(z-(x+\mu_{\phi^{*}}))^{T}K(z-(x+\mu_{\phi^{*}}))\}}$
$=\exp\{\frac{1}{2}(2\delta^{T}Kz-2\delta^{T}K(x+\mu_{\phi^{*}})-\delta^{T}K\delta)\}$
$=\exp\{\delta^{T}Kz-\delta^{T}K(x+\mu_{\phi^{*}})-\frac{1}{2}\delta^{T}K\delta\}$
$=\exp\{\delta^{T}Kz-\tau\}$
where $\tau=\delta^{T}K(x+\mu_{\phi^{*}})+\dfrac{1}{2}\delta^{T}K\delta$. This leads us to the following inequality:
$\exp\{\delta^{T}Kz-\tau\}\leq w\Longleftrightarrow\delta^{T}Kz\leq\log{w}+\tau=\eta.$
Leveraging the Neyman-Pearson theorem in [83], we establish that if the following condition holds for any binary classifier $\mathcal{F}_{\textrm{bin}}:\mathbb{R}^{d}\rightarrow\{0,1\}$:
$\text{Pr}(\mathcal{F}_{\textrm{bin}}(X)=1)\geq\text{Pr}(X\in\Omega)$
$\Omega=\{z\in$ $\mathbb{R}^{d}:\frac{p_{X^{*}}(z)}{p_{X}(z)}\leq w\}\textrm{ for some }w>0$
then it must hold that $\text{Pr}(\mathcal{F}_{\textrm{bin}}(X^{*})=1)\geq\text{Pr}(X^{*}\in\Omega)$. This statement is also satisfied when the direction of the inequality is reversed:
$\text{Pr}(\mathcal{F}_{\textrm{bin}}(X)=1)\leq\text{Pr}(X\in\Omega)$
$\Omega=\{z\in$ $\mathbb{R}^{d}:\frac{p_{X^{*}}(z)}{p_{X}(z)}\geq w\}\textrm{ for some }w>0$
indicating that $\text{Pr}(\mathcal{F}_{\textrm{bin}}(X^{*})=1)\leq\text{Pr}(X^{*}\in\Omega)$ is also true. This principle underpins our main claim. Suppose there exists a class $c_{A}$ within the set $\mathcal{C}$ with probability $p_{A}$ bounded by $\underline{p_{A}}$ and $\overline{p_{B}}$ such that:
$\text{Pr}(\mathcal{F}(x+\phi^{*})=c_{A})$ $\geq\underline{p_{A}}\geq\overline{p_{B}}\geq$
$\max_{c\neq c_{A}}\text{Pr}(\mathcal{F}(x+\phi^{*})=c).$
Then, it necessarily follows that:
$\mathcal{G}_{\phi^{*}}(x+\delta)$ $=c_{A},\;\forall\delta\in\{\delta:||\delta||_{2}<R\}$
where $\text{where }R$ $=\frac{1}{2}\sqrt{\lambda_{\min}}(\Phi^{-1}(\underline{p_{A}})-\Phi^{-1}(\overline{p_{B}})).$
The result means that our smoothed classifier $\mathcal{G}_{\phi^{*}}$ retains the most probable class of the original classifier even in the presence of input perturbation. Formally, it implies that $\text{Pr}(\mathcal{F}(X^{*})=c_{A})$ is at least as large as $\text{Pr}(\mathcal{F}(X^{*})=c)$ for any $c\in\mathcal{C}-\{c_{A}\}$. Without loss of generality, we fix $c$ as $c_{B}$ which is the second most probable class, then we need to show that $\text{Pr}(\mathcal{F}(X^{*})=c_{A})\geq\text{Pr}(\mathcal{F}(X^{*})=c_{B})$. The bounds for probabilities $\underline{p_{A}}$ and $\overline{p_{B}}$ can be expressed as $\text{Pr}(X\in\Omega_{A})$ and $\text{Pr}(X\in\Omega_{B})$ when the regions $\Omega_{A}$ and $\Omega_{B}$ are given by:
$\Omega_{A}=\{z:\delta^{T}K(z-(x+\mu_{\phi^{*}}))^{T}\leq\kappa\Phi^{-1}(\underline{p_{A}})\}$
$\Omega_{B}=\{z:\delta^{T}K(z-(x+\mu_{\phi^{*}}))^{T}\geq\kappa\Phi^{-1}(1-\overline{p_{B}})\}.$
In defining these regions, we denote $\kappa$ as the norm $=||\delta^{T}\Sigma_{\phi^{*}}^{-1/2}||_{2}.$ To further clarify the discussion, we define several useful random variables transformed from $X$:
$Y=X-(x+\mu_{\phi^{*}})\sim\mathcal{N}_{d}(0,\Sigma_{\phi^{*}})$
$Y^{*}=X-(x+\mu_{\phi^{*}}+\delta)\sim\mathcal{N}_{d}(\delta,\Sigma_{\phi^{*}})$
$\Xi=\Sigma_{\phi^{*}}^{-1/2}Y\sim\mathcal{N}_{d}(0,\mathbf{I})$
$Z\sim\mathcal{N}(0,1).$
We show $\text{Pr}(X\in\Omega_{A})=\underline{p_{A}}$ as follows:
$\text{Pr}(X\in\Omega_{A})$ $=\text{Pr}(\delta^{T}K(X-x-\mu_{\phi^{*}})\leq\kappa\Phi^{-1}(\underline{p_{A}}))$
$=\text{Pr}(\delta^{T}KY\leq\kappa\Phi^{-1}(\underline{p_{A}}))$
$=\text{Pr}(\delta^{T}K\Sigma_{\phi^{*}}^{1/2}\Xi\leq\kappa\Phi^{-1}(\underline{p_{A}}))$
$=\text{Pr}(\delta^{T}\Sigma_{\phi^{*}}^{-1/2}\Xi\leq\kappa\Phi^{-1}(\underline{p_{A}}))$
$=\text{Pr}(||\delta^{T}\Sigma_{\phi^{*}}^{-1/2}||_{2}Z\leq\kappa\Phi^{-1}(\underline{p_{A}}))$
$=\text{Pr}(Z\leq\Phi^{-1}(\underline{p_{A}}))$
$=\underline{p_{A}}.$
Here, $\delta^{T}\Sigma_{\phi^{*}}^{-1/2}$ is a row vector in $\mathbb{R}^{1\times d}$ and a random vector $\Xi$ in $\mathbb{R}^{d\times 1}$ follows a multivariate standard normal distribution. The inner product of theses vectors results in a weighted sum of univariate standard normal distributions, which reduces to a univariate normal distribution scaled by the constant $\kappa$. For $\Omega_{B}$, the probability is calculated as:
$\text{Pr}(X\in\Omega_{B})$ $=\text{Pr}(\delta^{T}K(X-x-\mu_{\phi^{*}})\geq\kappa\Phi^{-1}(1-\overline{p_{B}}))$
$=\text{Pr}(\delta^{T}KY\geq\kappa\Phi^{-1}(1-\overline{p_{B}}))$
$=\text{Pr}(\delta^{T}K\Sigma_{\phi^{*}}^{1/2}\Xi\geq\kappa\Phi^{-1}(1-\overline{p_{B}}))$
$=\text{Pr}(\delta^{T}\Sigma_{\phi^{*}}^{-1/2}\Xi\geq\kappa\Phi^{-1}(1-\overline{p_{B}}))$
$=\text{Pr}(||\delta^{T}\Sigma_{\phi^{*}}^{-1/2}||_{2}Z\geq\kappa\Phi^{-1}(1-\overline{p_{B}}))$
$=\text{Pr}(Z\geq\Phi^{-1}(1-\overline{p_{B}}))$
$=\overline{p_{B}}.$
We can modify our initial assumption by replacing $\underline{p_{A}}$ to $\text{Pr}(X\in\Omega_{A})$ and $\overline{p_{B}}$ to $\text{Pr}(X\in\Omega_{B})$.
$\text{Pr}(\mathcal{F}(X)=c_{A})\geq\text{Pr}(X\in\Omega_{A})$
$\text{Pr}(\mathcal{F}(X)=c_{B})\leq\text{Pr}(X\in\Omega_{B}).$
We utilize Neyman-Pearson theorem by defining $\mathcal{F}_{\textrm{bin}}(z)=I[\mathcal{F}(z)=c_{A}]$ where $I$ is an indicator function, then the result are ensured
$\text{Pr}(\mathcal{F}(X^{*})=c_{A})\geq\text{Pr}(X^{*}\in\Omega_{A})$
$\text{Pr}(\mathcal{F}(X^{*})=c_{B})\leq\text{Pr}(X^{*}\in\Omega_{B}).$
If $\text{Pr}(X^{*}\in\Omega_{A})\geq\text{Pr}(X^{*}\in\Omega_{B})$ is satisfied, $\text{Pr}(\mathcal{F}(X^{*})=c_{A})\geq\text{Pr}(\mathcal{F}(X^{*})=c_{B})$ is held with a certain radius, which is the destination of our proof.
$\text{Pr}(\mathcal{F}(x+\phi+\delta)=c_{A})\geq\max_{c\neq c_{A}}\text{Pr}(\mathcal{F}(x+\phi+\delta)=c).$
To show the inequality, we plug in the space $\Omega_{A}$ we defined into $\text{Pr}(X^{*}\in\Omega_{A})$
$\text{Pr}(X^{*}\in\Omega_{A})$
$=\text{Pr}(\delta^{T}K(X^{*}-x-\mu_{\phi^{*}})\leq\kappa\Phi^{-1}(\underline{p_{A}}))$
$=\text{Pr}(\delta^{T}KY^{*}\leq\kappa\Phi^{-1}(\underline{p_{A}}))$
$=\text{Pr}(\delta^{T}KY+\delta^{T}K\delta\leq\kappa\Phi^{-1}(\underline{p_{A}}))$
$=\text{Pr}(\delta^{T}\Sigma_{\phi^{*}}^{-1/2}\Xi\leq\kappa\Phi^{-1}(\underline{p_{A}})-\delta^{T}K\delta)$
$=\text{Pr}(\kappa Z\leq\kappa(\Phi^{-1}(\underline{p_{A}})-\frac{1}{\kappa}\delta^{T}K\delta))$
$=\text{Pr}(Z\leq\Phi^{-1}(\underline{p_{A}})-\frac{1}{\kappa}\delta^{T}K\delta)$
$=\Phi(\Phi^{-1}(\underline{p_{A}})-\frac{1}{\kappa}\delta^{T}K\delta).$
We follow a similar procedure for the space $\Omega_{B}$
$\text{Pr}(X^{*}\in\Omega_{B})$
$=\text{Pr}(\delta^{T}K(X^{*}-x-\mu_{\phi^{*}})\geq\kappa\Phi^{-1}(1-\overline{p_{B}}))$
$=\text{Pr}(\delta^{T}KY^{*}\geq\kappa\Phi^{-1}(1-\overline{p_{B}}))$
$=\text{Pr}(\delta^{T}KY+\delta^{T}K\delta\geq\kappa\Phi^{-1}(1-\overline{p_{B}}))$
$=\text{Pr}(\delta^{T}\Sigma_{\phi^{*}}^{-1/2}\Xi\geq\kappa\Phi^{-1}(1-\overline{p_{B}})-\delta^{T}K\delta)$
$=\text{Pr}(\kappa Z\geq\kappa(\Phi^{-1}(1-\overline{p_{B}})-\frac{1}{\kappa}\delta^{T}K\delta))$
$=\text{Pr}(Z\geq\Phi^{-1}(1-\overline{p_{B}})-\frac{1}{\kappa}\delta^{T}K\delta)$
$=\text{Pr}(Z\leq\Phi^{-1}(\overline{p_{B}})+\frac{1}{\kappa}\delta^{T}K\delta)$
$=\Phi(\Phi^{-1}(\overline{p_{B}})+\frac{1}{\kappa}\delta^{T}K\delta).$
By comparing these two probabilities, we identify the condition under which $\text{Pr}(X^{*}\in\Omega_{A})\geq\text{Pr}(X^{*}\in\Omega_{B})$ holds as follows:
$\Phi(\Phi^{-1}(\underline{p_{A}})-\frac{1}{\kappa}\delta^{T}K\delta)\geq\Phi(\Phi^{-1}(\overline{p_{B}})+\frac{1}{\kappa}\delta^{T}K\delta)$
$\Longleftrightarrow\quad\Phi^{-1}(\underline{p_{A}})-\frac{1}{\kappa}\delta^{T}K\delta\geq\Phi^{-1}(\overline{p_{B}})+\frac{1}{\kappa}\delta^{T}K\delta$
$\Longleftrightarrow\quad\frac{\delta^{T}K\delta}{\kappa}\leq\frac{1}{2}(\Phi^{-1}(\underline{p_{A}})-\Phi^{-1}(\overline{p_{B}})).$
$\kappa$ is previously defined as $||\delta^{T}\Sigma^{-1/2}_{\phi^{*}}||_{2}$, which is $\sqrt{\delta^{T}\Sigma^{-1}_{\phi^{*}}\delta}$. Since $\Sigma_{\phi^{*}}$ is a Hermitian matrix, $K=\Sigma^{-1}_{\phi^{*}}$ is also a Hermitian matrix, satisfying the condition for Rayleigh quotient in [87]. The upper bound of $\frac{\delta^{T}K\delta}{\kappa}$ is $\frac{||\delta||}{\sqrt{\lambda_{\min}}}$, as determined by Rayleigh quotient:
$\frac{\delta^{T}K\delta}{\kappa}=\frac{\delta^{T}\Sigma^{-1}_{\phi^{*}}\delta}{\sqrt{\delta^{T}\Sigma^{-1}_{\phi^{*}}\delta}}=\sqrt{\delta^{T}\Sigma^{-1}_{\phi^{*}}\delta}\leq\sqrt{\frac{||\delta||^{2}}{\lambda_{\min}}}=\frac{||\delta||}{\sqrt{\lambda_{\min}}}.$
Here, $\lambda_{\min}$ stands for the minimum eigenvalue of $\Sigma_{\phi^{*}}$. The right side is derived by the equality $\lambda_{\max}(\Sigma_{\phi^{*}}^{-1})=\frac{1}{\lambda_{\min}(\Sigma_{\phi^{*}})}$, where $\lambda_{\max}(\Sigma_{\phi^{*}}^{-1})$ denotes the maximum eigenvalue of the matrix $\Sigma^{-1}_{\phi^{*}}$, and $\lambda_{\min}(\Sigma_{\phi^{*}})$ is the minimum eigenvalue of $\Sigma_{\phi^{*}}$ denoted as $\lambda_{\min}$. When the upper bound is less than or equal to $\frac{\kappa}{2}\Psi$ such that $\Psi=\Phi^{-1}(\underline{p_{A}})-\Phi^{-1}(\overline{p_{B}})$, $\mathcal{G}_{\phi^{*}}(x+\delta)=c_{A}$ is supported.
$\frac{||\delta||}{\sqrt{\lambda_{\min}}}\leq\frac{1}{2}\Psi\Longleftrightarrow||\delta||\leq\frac{1}{2}\sqrt{\lambda_{\min}}\Psi=R.$
In other words, our classifier with Gaussian visual prompt ensures the robustness against the perturbation $\delta$ by $R$. Also, for any vector $x$ in $\mathbb{R}^{d}$, the upper bound of $\lambda_{\min}$ is introduced by Rayleigh quotient.
$\lambda_{\min}$ $\leq\min_{x\neq 0}{\frac{x^{T}\Sigma_{\phi^{*}}x}{x^{T}x}}$
$=\min_{\sum^{d}_{i=1}x_{i}^{2}=1}{(x_{1}^{2}\sigma_{1}^{2}+\cdots+x_{d}^{2}\sigma_{d}^{2})}$
$=\min\{\sigma_{i}^{2}\}^{d}_{i=1}.$
This induces the upper bound of $R$, which is $\overline{R}=\frac{1}{2}\min\{\sigma_{i}\}^{d}_{i=1}\Psi$.
Appendix B Experimental Setting
B-A Datasets
Synthetic Datasets
Our BlackVIP generates the input-dependent image-size visual prompt which covers the whole image region, so we expect that this flexible prompt design can improve some kind of robustness as well as general recognition capability: (1) To evaluate the robustness on distribution shift (i.e., domain generalization), we consider the Biased-MNIST [70] dataset. (2) To evaluate the robustness on adversarial noise and location-agnostic recognition capacity, we create a variant of the MNIST dataset called Loc-MNIST. Examples of these two datasets are provided in Figure 10.
Figure 8: Examples of $y=7$ subset in Biased-MNIST [70] with $\rho=0.9$. (Top) The train set is constructed with the spurious correlation between the background color and digit class (e.g., $y=7$ occurs $90\%$ with a pink background and $10\%$ with other random colors in this case). (Bottom) The test set is constructed with a reversed correlation to that of the train set (e.g., $y=7$ occurs $10\%$ with a pink background and $90\%$ with other random colors in this case).
Figure 9: Examples of the Loc-MNIST dataset. The real digit from MNIST is located in the outer area, while the fake digit from another random MNIST image is placed in the center of the image. (Top) The case where the size ratio of the real digit to the fake digit is 1:1, and (Bottom) 1:4.
Figure 10: Examples of two synthetic datasets. (a) Biased MNIST and (b) Loc-MNIST
Biased MNIST is a modified version of MNIST [71] where the biases reside in the background colors of the images of each digit. At train time, each digit has a unique preassigned background color that strongly correlates with the label. The degree of correlation is determined by the value $\rho\in[0,1]$, such that $(100\times\rho)\%$ of the images that belong to the same digit have the preassigned color of that digit as their background color, and the rest are uniformly assigned to have any of the other colors as their background color. At test time, we reverse the ratio so that $(100\times(1-\rho))\%$ of the images now have the preassigned color as their background color and vice versa to evaluate the model’s dependency on superficial features such as the color of the background that a digit is located on. We prepare the following two environments 1) easy: $\rho=0.8$ and 2) hard: $\rho=0.9$.
On the given black blank image with $224\times 224$ resolution, i.e., zero array, Loc-MNIST puts an original target digit image from MNIST that has $28\times 28$ resolution on the edge-side (e.g., 0$\sim$27 or 196$\sim$223 for one of the vertical or horizontal sides and 0$\sim$223 for the other side) and puts a random fake digit (also from the MNIST dataset) on the center. The location of the target digit in the edge and the class of the fake digit are chosen randomly with uniform probability. A synthetic image is created one by one for each original MNIST image. We prepare the following two environments 1) easy: the scale of the target and the fake digit is the same, i.e., 1:1, and 2) hard: the fake digit is four times larger than the original digit, i.e., 1:4.
For consistency, we perform the experiments on these two datasets with a few-shot evaluation protocol. To construct a train set, we randomly sample a subset (K-shot) of the created images for each class and use the whole test set.
Datasets
To extensively evaluate the effectiveness of our proposed method and baseline approaches, we measure performance across the following 14 datasets that are widely used for transfer learning benchmark: Caltech101 [88], OxfordPets [89], StanfordCars [90], Flowers102 [91], Food101 [92], FGVCAircraft [93], SUN397 [94], DTD [72], SVHN [95], EuroSAT [73], Resisc45 [74], CLEVR [75], UCF101 [96], and ImageNet (IN) [97]. Note that these 14 datasets cover diverse visual domains, and they require understanding various visual semantics like scenes, actions, fine-grained categories, textures, satellite imagery, digits, the number of objects, and the recognition of generic objects. Moreover, to evaluate the robustness of prompt-based adaptation of PTM in the wild, we additionally consider three benchmark datasets of distribution shifts from WILDS [68]: Camelyon17 [98], FMoW [99], and iWildCam [100].
Following the protocol in [10, 11], we conduct a few-shot evaluation for all datasets: 16-shot for the train set, 4-shot for the validation set, and the whole test set. We use the few-shot split by [10] for each dataset, those are also used in [10], while for Resisc45, CLEVR, and WILDS datasets, we randomly select the 16-shot and 4-shot samples for the training and validation datasets, respectively.
B-B Backbone Model
In this work, we aim at the robust adaptation of pre-trained models on diverse downstream tasks. For these pre-trained models, all experiments in this paper are done with the off-the-shelf vision-language model CLIP [1], and we adopt the ViT-B/16 for image encoder backbone architecture by default. During the adaptation (training) phase, the entire components of the pre-trained model are frozen without any architectural modification, and we only manage and optimize the learnable module Coordinator from the outside of the pre-trained model.
While input space visual prompting allows it to be applied to not only VLM, but also any other vision models like CNNs and ViTs, it requires the user to define the output space mapping, which maps the output prediction category set of a pre-trained task to a new downstream category set [40, 31, 6]. This is another non-trivial problem. Therefore, we limit our focus to only the VLM that can dynamically build the task-specific head from a manual text template [1, 101] so that free from defining output space mapping.
B-C Baseline Methods
CLIP Zero-Shot (ZS)
CLIP [1] is one of the most popular vision-language zero-shot models that is widely exploited for classification, detection, segmentation, and other vision or vision-language tasks. Based on its well-aligned vision-language joint embedding space, the zero-shot classification can be performed with a manual text prompt (also called a template) of each pre-defined class category. In this paper, we are mainly aiming to improve CLIP’s strong zero-shot performance in the few-shot adaptation setting.
BAR
Black-Box Adversarial Reprogramming (BAR) [31] was proposed for efficient transfer learning of a pre-trained model to the medical image domain. Different from the previous works on Adversarial Reprogramming (AR), BAR exploits the perturbation-vulnerability of neural networks for adaptation purposes rather than attack. By optimizing the frame-shaped learnable program, which embeds a downstream target image inside of that, BAR steers the ImageNet pre-trained model to classify the specialized medical images. Moreover, BAR adopts the zeroth-order optimizer (ZOO), Randomized Gradient-Free (RGF) [102] minimization algorithm for black-box transfer learning to broaden its applications.
When the resolution of the downstream input image is over that of the pre-training phase, Tsai et al. [31] set the embedded target image size for $64\times 64$ resolution in the $299\times 299$-size learnable program by default. However, we observe that such a heavy-pad thin-image design of the prompt degrade the performance significantly, so we tune the resolution of the embedded image and set $194\times 194$.
VP
Similarly, Visual Prompting (VP) aims at adapting a pre-trained model to downstream tasks via learning input space visual prompts. Among some candidates for prompt designs, Bahng et al. [6] adopt the padding-style prompt so that realized prompts look like the frame-shape program of ARs. VP learns a universal visual prompt per each downstream task, and it just adds to all of the images in a task. Unlike the AR methods or our BlackVIP, the range of prompted images is unbounded. Following [6], we use the padding-style prompt, which is 30-pixel sized for each side by default.
While VP optimizes the parameters in the input space, it relies on a first-order optimization algorithm that uses the true gradient of the entire model parameters, and we establish the performance of VP as a kind of upper bound for other input space black-box optimization approaches, including BlackVIP and BlackVIP-SE. Additionally, by replacing the first-order algorithm with zeroth-order counterparts, we build two new baselines VP w/ SPSA and VP w/ SPSA-GC on our extensive experiments. These two methods confirm the effectiveness of our new components Coordinator and SPSA-GC.
Furthermore, we consider some representative prompt learning methods as baselines on language modality (CoOp [22] and CoCoOp [11]), vision modality (VP [6] and VPT [5]), and both modality (MaPLe [7]) for an evaluation of white-box transfer learning experiments to validate the effectiveness of the design of our visual prompt, solely. All of these methods adapt parameters inside the model, e.g., tokens in the embedding layer or intermediate layers of text or visual encoder of the target vision-language model.
Discussion
Although BAR, VP, and BlackVIP share the generic goal: efficient transfer learning of pre-trained models via input-space optimization, there are several significant differences. (1) We propose a novel prompt design that is automatically formed in an input-dependent manner rather than the frame-shaped manual design of the input-independent prompt (or program) of VP (or BAR). (2) While VP relies on first-order algorithms and BAR adopts the RGF, we utilize the new variants of SPSA [39], SPSA-GC, which is enhanced with a proper modification in the parameter update rule. (3) Contrary to the medical imaging-only validation in BAR, based on the above two technical difference, BlackVIP successfully adapt the pre-trained model to diverse data domains (described in Section B.1.).
B-D Implementation Details
Architecture
For the fixed text prompt design of each dataset, those are shared across all baseline methods and BlackVIPs, we use the same templates provided by [6] for SVHN, CLEVR, and Resisc45, and [10] for the remaining 11 datasets. For the fixed encoder part of our Coodinator, we use the ImageNet pre-trained vit-mae-base checkpoint [^3] https://huggingface.co/docs/transformers/model_doc/vit_mae from HuggingFace. The output shape of the encoder is $N\times 768$, where $N$ is the number of instances in the batch. For the BlackVIP-SE, we used scikit-learn’s vanilla PCA or kernel PCA method [^4] We considered whether to use a kernel or not and which kernel (among RBF, cosine, and polynomial kernels) to use as hyperparameters.. For fast training, we pre-computed the PCA features for BlackVIP-SE and retrieved them by index during each training iteration. During the inference phase, we load the trainset-fitted PCA projection matrix and apply it to the flattened test input $x$ to get the feature vector fed into the prompt decoder. Here, we design the decoder based on depth-wise separable convolution (DSC) layer [103] for parameter efficiency. Specifically, we build a block of [NORM-ACT-CONV] and stack it five times. The NORM and ACT denote Batch Normalization and Gaussian Error Linear Unit, respectively. The CONV operation of the first four blocks is DSC, and the last one is a standard convolutional layer. Our implementation code is available at https://github.com/changdaeoh/BlackVIP
To satisfy a fully convolutional design without loss of expressiveness, tensors that are fed into the decoder must be shaped in a 3D feature map. For this, we additionally govern a task-specific single continuous vector $\phi_{t}$ (called prompt trigger vector), which is concatenated with the output feature vector of encoder, leading to the appropriate size of 1D vector for reshaping to a 3D tensor. For the BlackVIP, we set the dimension of the prompt trigger vector to 800, resulting in 1568 dimensions of a concatenated vector that can be reshaped to $32\times 7\times 7$ shaped 3D tensor. Meanwhile, BlackVIP-SE allows users to set the latent feature dimension (for the reduced PCA coordinate system) according to their own preference. This contributes to further reducing the input dimension of the decoder feature map, and we set the PCA feature dimension to 98. Rather than concatenation, we simply use the sum of the trigger vector and PCA feature as input to the prompt decoder. The prompt trigger is shared across all instances for a given task.
Optimization and other configurations
For a stable approximation of the gradient in practice, ZOO algorithms repeat the gradient estimation step for several times and use the mean of those estimates as a final approximation of the gradient. Usually, the approximation quality is proportional to the number of these repeats. We set this repeat as five times for all baselines that use ZOO.
Besides the learning rate and learning rate schedule parameters, ZOO algorithms have some additional algorithm-specific hyperparameters that need to be tuned. For RGF, these are the standard deviation of a random Gaussian vector and a smoothing parameter, and for SPSA, these are the perturbation magnitude and its decaying factor. We provide the search range of each hyperparameter in Table VIII. The search range for algorithm-specific parameters is based on the proposal of authors of SPSA [13] and BAR [31]. Moreover, among the valid perturbation distributions of SPSA, we adopt the Segmented Uniform $[-1.0,-0.5]\cup[0.5,1.0]$.
The learning objective is a cross-entropy loss for VP and BlackVIP and a focal loss for BAR (following [31]). For all black-box approaches, the batch size is set to 128 across all datasets. Except for the SUN397 (1,000), StanfordCars (2,500), and ImageNet (500), we optimize all methods during 5,000 epochs for convergence. Note that the input space visual prompting with first-order algorithm already requires sufficiently large iterations, e.g., 1,000 epochs [6] with full dataset, and ZOO demands many more iterations due to the lack of gradient information.
B-E Hyperparameter Sweep
In this section, we provide the hyperparameter search range of each algorithm, summarized in Table VIII.
TABLE VIII: Hyperparameter sweep. Large LR (learning rate) of BAR and VP is based on [6] to directly optimize pixel values rather than the neural network’s weights. PM denotes perturbation scale $c_{i}$, and parameters of the dimensionality reduction algorithm.
Hyperparameter Algorithm Search Range
[…]
Appendix C Detail Description of BlackVIP’s Coodinator
On the transfer learning of a pre-trained model which provides no accessibility about any architectural information or actual model parameters, BlackVIP treats this situation with two novel mechanisms: (1) parameter-efficient instance-aware prompt generation network, and (2) stable zeroth-order optimization algorithm that is based on SPSA [13]. In this section, we provide a detailed description of the first component, Coordinator.
Different from existing works on visual prompting, we reparameterize the input space visual prompt $\phi$ as a neural network, Coordinator $v_{\bm{\phi}}(\cdot)$ that generates an input-dependent visual prompt $v_{\bm{\phi}}(x)$. Coordinator is composed with encoder $f(\cdot)$, decoder $g_{\phi_{d}}(\cdot)$ and task-specific learnable vector $\phi_{t}$. The encoder is used for extracting instance-specific latent feature vector $z_{x}=f(x)$ contributing to the construction of the optimal input space visual prompt for each instance. Because our goal in this work is the broad utilization of pre-trained models on diverse downstream tasks, we adopt a pre-trained encoder network optimized by a self-supervised learning objective, not by a supervised learning objective or scratch network. Specifically, we use the ViT-B/16 weights from the Masked AutoEncoding pre-training [48]. We present the grounds for using the self-supervised learning encoder in the main paper, refer to Sec. 3. During the training phase, this pre-trained encoder part is frozen (not updated) and just acts as a feature extractor. Then, the instance-specific feature vector from the encoder is conveyed to the decoder for a prompt generation.
Prompt decoder $g_{\phi_{d}}(\cdot)$ is a lightweight convolutional neural network, which has learnable parameters less than 10K by default. Note that the generated prompt has the same shape as the input image, so our prompt covers the entire region of the image, unlike previous visual prompting and reprogramming works applied to the partial region of the image.
In addition to the feature vector from the fixed encoder, the decoder also incorporates an additional input which is shared for all instances across the current dataset. The so-called prompt trigger vector $\phi_{t}$ is a continuous vector that also contributes to the design of a visual prompt by collaborating with the instance-specific rich feature vector from the encoder. By introducing this prompt trigger vector, the decoder of the Coordinator can enjoy additional information to generate more proper prompts for a given task. Besides, it helps to build the 3D feature map for the decoder’s input, which is necessary for designing a parameter-efficient fully convolutional decoder network.
Appendix D Additional Experimental Results
D-A Further Analysis on Synthetic Dataset
We validated our novel prompt design strategy on some challenging generalization setups, i.e., correlation shift (Biased MNIST) and varying object location (Loc-MNIST). As supporting evidence for the improved performance on these sets, we provide the visualization of the t-SNE embedding and normalized mutual information between the visual embeddings of the final layer and the ground truth labels in Fig. 11, where we can see that BlackVIP and BlackVIP-SE successfully discriminate different classes of images in the input space, resulting in better discriminative latent embeddings.
Figure 11: (Left) Embedding visualization with t-SNE [104] on the prompt-augmented images from test splits of Loc-MNIST 1:4 setup. (right) Normalized Mutual Information (NMI) [105] score of the learned embedding.
D-B Ablation Study
D-B1 Model Architectures
To study the versatility of our method, we vary the backbone architecture of the pre-trained target model and the encoder of Coordinator in Tab. IX. While BAR and the naive application of SPSA-GC on VP fail to improve the zero-shot performance of CNN-based target backbones that lack the global attention of Transformers [21], our BlackVIPs consistently bring huge performance gains across all the architectures. It implies that BlackVIPs are architecture-agnostic approaches, which pursue the general adaptation method for high-performing PTMs. Regarding the comparison between BlackVIP and BlackVIP-SE, BlackVIP shows better results when the model architecture of the Coordinator encoder aligns with the target PTM’s architecture. Meanwhile, BlackVIP-SE achieves more stable performance improvement across all architectures.
TABLE IX: Ablation study for backbone architecture. Classification accuracy on EuroSAT across pre-trained target backbone architectures and BlackVIP’s Coordinators.
Method Target Backbone
RN50 RN101 ViT-B/32 ViT-B/16 Avg.
ZS 37.5 32.6 45.2 40.8 48.4
BAR 26.9 33.5 70.3 77.2 52.0
VP w SPSA-GC 34.7 31.2 71.1 70.9 52.0
BlackVIP-SE 50.0 55.1 66.2 71.2 60.6
BlackVIP (RN50) 51.3 50.8 62.9 68.5 58.4
BlackVIP (ViT-B/16) 48.4 51.3 67.9 73.1 60.2
TABLE X: Different Coordinator weights with SPSA variants. Mean classification accuracy of three repeated runs on EuroSAT.
Encoder Type Optim. Acc.
Zero-Shot 47.9
scratch SPSA 49.6
scratch SPSA-GC 49.5
Sup. pre-trained SPSA 59.4
Sup. pre-trained SPSA-GC 65.2
SSL pre-tained SPSA 69.4
BlackVIP-SE (PCA matrix with SPSA-GC) 71.2
BlackVIP (SSL pre-trained with SPSA-GC) 73.1
D-B2 Coordinator and Zeroth-order Methods
BlackVIPs adopt the encoder-decoder structure to efficiently generate the input-dependent prompts. We exploit an SSL pre-trained encoder (in BlackVIP) and PCA projection encoder (in BlackVIP-SE) while learning a randomly initialized lightweight decoder from scratch. From the design philosophy of BlackVIPs, we expect that the encoder extracts the essential features of the given image, including the spatial features, and the decoder utilizes the features to produce a spatially and semantically structured prompt tailored to the input. We conjecture that SSL pre-trained and PCA encoders are desirable to capture the demanding semantics instead of a supervised one learned from pre-defined labels. Here, Table X confirms that the SSL encoder outperforms the supervised pre-trained or randomly initialized encoder (scratch). Besides, the PCA encoder also outperforms the large-scale supervised encoder, which indicates the potential of the classic statistical methods for feature extraction during transfer learning. Furthermore, SPSA-GC improves the 3.7% accuracy than SPSA, from 69.4 to 73.1. It denotes that approximated gradients by our SPSA-GC are more accurate than the original SPSA.
D-C Query Efficiency
Figure 12: Classification accuracy for given queries and corresponding budget (\$ USD) of different black-box visual prompting method.
We present an expanded analysis of the query efficiency experiment in the manuscript on four model datasets covering the natural object recognition (Pets), action recognition (UCF101), and fine-granular visual recognition (Cars and Aircraft). We observe outstanding query efficiency of BlackVIP and BlackVIP-SE compared with other black-box visual prompting methods, which commonly failed to achieve meaningful improvement from zero-shot inference under a few queries.
D-D Visual Prompting in the Wilds
TABLE XI: Robustness comparison under real-world distribution shift.
Method iWildCam FMoW Camelyon17 Avg.
ZS 9.9 18.9 55.2 28.0
BAR 11.0 22.4 76.8 36.7
VP w/ SPSA-GC 10.9 22.3 66.5 32.5
BlackVIP-SE 10.6 21.9 79.7 37.4
BlackVIP 12.3 22.6 76.5 37.1
These days, our AI models commonly encounter distribution shifts in real-world applications. In the manuscript, we evaluate the robustness of prompting-based adaptation approaches against synthetic correlation shift environments. We now validate those approaches on three kinds of real-world datasets (iWildCam, FMoW, and Camelyon17) from the WILDS distribution shift benchmark [68]. In iWildCam, the task is to classify the 182 species of animals in the wild under different remote sensing conditions (location of camera traps). In FMoW, satellite images should be classified into the building and land types, 62 in total, and the shifting factor is the combination of (year, region). Meanwhile, Camelyon17 requires a binary classification (tumor or normal) on the patch-level tissue slides under varying hospital imaging conditions. In Table XI, the baseline approaches, BAR and VP w/ SPSA-GC, show their effectiveness in improving the zero-shot classifier’s performance, and our BlackVIPs achieve superior average performances. Interestingly, while BlackVIP-SE underperforms other methods in iWildCam and FMoW, it shows outstanding performance on Camelyon17, where we speculate the intrinsic dimensionality of the task is relatively low, and the PCA features can be discriminative enough to generate good prompts.
D-E Grad-CAM Analysis
To investigate whether visual prompts produced by each method adapt the pre-trained model, we visualize the Grad-CAM [106] on the original image and prompted image of the encoder’s penultimate layer (Figure 13-20). We select eight datasets that represent the diverse image domains and experimental conditions: (Natural) OxfordPets and SVHN, (Specialized) EuroSAT, (Structured) CLEVR-count, (Action Recognition) UCF101, (Fine-Grained) StanfordCars, (Synthetic) Biased-MNIST and Loc-MNIST. Detail descriptions for each dataset are provided in Sec. A.1.
BlackVIP generates diverse prompts to properly adapt the pre-trained model to the targeted data domains. When the task is counting the number of objects (CLEVR) in the entire region of an image, BlackVIP extends the attention of the pre-trained model to the whole view of an image as shown in Figure 13. If the task requires a fine-grained understanding of objects or recognition of actions (UCF101), BlackVIP concentrates the model’s attention on class-related regions, as shown in Figure 14.
Figure 13: Grad-CAM on CLEVR. Compared to baseline methods, BlackVIP extends the attention of models to broad areas of the image for effective reasoning on the number of objects.
Figure 14: Grad-CAM on UCF101. Compared to baseline methods, BlackVIP concentrates the attention of models on local areas of the image for effective recognition of the specific actions.
Figure 15: Grad-CAM on OxfordPets. Compared to baseline methods, BlackVIP effectively adapts the model to focus on the target object rather than spurious features such as the background.
Figure 16: Grad-CAM on SVHN. Compared to baseline methods, BlackVIP effectively adapts the model to focus on the target digit rather than spurious features such as the background.
Figure 17: Grad-CAM on EuroSAT. Compared to baseline methods, BlackVIP extends the attention of models to broad areas of the image for effective classification of satellite imagery.
Figure 18: Grad-CAM on StanfordCars. Compared to baseline methods, BlackVIP concentrates the attention of models on an object or local areas of an image for effective fine-grained classification.
Figure 19: Grad-CAM on Biased-MNIST. While baseline methods attend to the background rather than digit shape, our BlackVIP can bypass this spurious feature through a widely scattered visual prompt and focus more of the attention on the shape of the digit.
Figure 20: Grad-CAM on Loc-MNIST. Compared to baseline methods, BlackVIP effectively adapts the model to aim at edge-located true digit corresponding true label rather than the obstructive fake digit in the center of the image.